# REACT: Real-time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation

### v2

@@ B Model mR@K R@K F1@K mAP50
SQUAT [16] 15.3 26.7 19.4 -
BGNN [19] 11.6 33.4 17.3 29.0
@@
OS EGTR [10] 7.8 29.3 12.4 30.8
SGTR [20] 13.6 26.5 18.0 25.4
SS-RCNN [37] 9.5 36.0 15.0 23.8
RelTR [4] 8.0 24.3 12.0 26.4
Relation Prediction Accuracy. On the PSG dataset, we experienced an average improvement of 58.47% in F1@k by using our DTS method compared to the same approaches with Faster-RCNN. In addition, our REACT model outperforms the PE-NET model with DTS by a slow margin (23.9 versus 23.2 in F1@K). We observe that the gain in F1@k is not very strongly correlated to the gain in mAP between the Two-Stage and DTS approaches, as we measured a correlation coefficient of 0.75 between the two. On the IndoorVG dataset, we observe a relative improvement of 38.76% in F1@K with the DTS approach. Our REACT model also outperforms PE-NET (DTS) by a strong margin (22.8 versus 20.7 in F1@K). In Table 4 we compare the performance of the REACT model on the VG150 dataset. We observe that REACT outperforms some Two-Stage approaches (Motifs, VCTree, and BGNN) but is not competitive with more recent approaches (PE-NET and SQUAT). We believe this is due to noisy data as well as the poor performance in OD of the YOLO backbone.
@@
Latency and Parameter Count. Regarding latency, we demonstrated an improvement of 84.99% using YOLOV8 instead of Faster-RCNN, which is a considerable gap. Regarding the number of parameters, the DTS approach uses on average 77.47% less parameters than the baseline. Our introduced REACT model is the fastest one with a latency of 32.5ms. By applying our DCS strategy, we gain an average of 66.5% latency on all models (compared to the YOLOV8m-DTS versions) with a relative average loss of 1% in F1@K. The EGTR model stays the best overall in terms of parameter count by a slow margin (42.5M versus 43.3M for REACT). By comparing REACT to the fastest model benchmarked, RelTR [4], REACT is 2.7x faster with $+10$ points on average in F1@K and $+24.9$ points on average in mAP50 for the 3 datasets. By taking into account OD accuracy, RelPred accuracy, and latency, our REACT model is the best compromise. Even for the VG150 dataset, REACT is a better choice because no other two-stage approaches are close to real-time inference. On the other hand, when evaluated fairly, one-stage approaches are clearly behind in terms of mAP and F1@K on all datasets, and are then not a good choice in comparison to REACT.
6 Ablation Studies
@@
Latency and Number of Proposals. As explained before, the second stage of SGG takes as input a fixed number of $n$ proposals as node candidates. There could be a maximum of $n*(n-1)$ possible pairs in the graph, thus when doing matching the computational complexity is supposed to scale accordingly. To demonstrate this hypothesis, we evaluated the performance of the REACT model in both latency and accuracy for different numbers of input proposals $k=\{10,150\}$ with DTS-YOLOV8 and with classical Faster-RCNN. For all experiments, we ranked proposals by confidence of the backbone and selected the top $k$.
@@
Figure 3: Top: Latency for the REACT model using a different number of proposals per image, with batch size 1. Bottom: Average F1@k for the REACT model with DTS-YOLO.
@@
The results of these experiments are shown in Fig. 3 (top), where we observe a plateau phenomenon for REACT with Faster-RCNN above 100 proposals. This behavior does not appear when using DTS+YOLOV8. By using 150 proposals with DTS+YOLOV8, the latency was similar to 10 proposals with Faster-RCNN. The evaluation of the REACT model under different sampling conditions with our DCS strategy is represented in Fig. 3 (bottom). In this figure, we can see that the optimal value was found at 23.89 for $k=42$, just before that the performance starts to be saturated.
Figure 4: Average Recall@K, meanRecall@K, and mAP@50 performance for the REACT model with DCS against the corresponding latency using different variants of the YOLOV8 model.
Scaling Down YOLOV8. To better evaluate the impact of the object detector accuracy on the relation prediction stage, we ran a set of experiments using different variants of YOLOV8 with the REACT model for Scene Graph Generation. Here, we trained the model with the same hyperparameters as before, only modifying the input size of the visual features used for relation prediction to match the different output channel sizes for each YOLOV8 variant. The results are displayed in Figure 4. We observed a strong correlation between mAP@50 and R@K (correlation coefficient = 0.97), and a correlation between mAP@50 and mR@K (correlation coefficient = 0.86). The performance of the larger model, YOLOV8-L is slightly worse in mR@K than the medium YOLOV8-M. We believe that the higher number of dimensions in the features extracted from the backbone could lead to more difficulties for the relation prediction training to converge. In addition, we observed a linear increase in latency with the size of the model. The performance of REACT with the smaller model, YOLOV8-N, is equivalent to the original PE-NET model on the PSG dataset but 22x faster and with 32x fewer parameters (426.5M versus 13.25M). In terms of the overall performance, the medium version seems to be the best. It has a good trade-off between all metrics and latency, making it a good choice for high performance in edge settings.
@@
Features Fusion. During context learning of the REACT model, and more generally in a lot of approaches in SGG [35, 44, 26, 46], different feature sources are used to model the edge representation. First, union and spatial features are combined and then merged with the addition of the subject and object node features. The node features are created using visual and textual features obtained for each proposal. In our REACT model, we removed the dependency on the union features (see Figure 2(d) (i)). In the following, we justify our choice by analyzing the impact of each feature type on learning. The results from these experiments are displayed in Table 5. The baseline REACT model is the second row. With union features (first row) we observe a significant decrease in F1@K as well as latency. We hypothesize that union features are too noisy and confuse the model. This is consistent with previous work which highlights the importance of spatial features for the task [12]. In the last two rows, we removed all dependencies on contextual features (union and spatial) and kept only the concatenation of the subject and object node as initial edge features. We observe only a slight decrease by keeping only text features (last row). In fact, the model can retain up to 71% of the optimal performance by learning only from the textual GLoVe features. This is an interesting finding which questions the balance between text and visual features in the task.
Table 5: Feature sources experiments. V and T are visual and textual features for $<subject,object>$, and S and U are Spatial and Union features, respectively.
@@
Feature Size. Previous two-stage approaches such as PE-NET use a fixed size of 4096 for both the edge and node representation. However, this design was not justified in previous work. Here, we ran experiments to benchmark the efficiency of different feature sizes for learning. We trained the REACT model with four different sizes, see Table 6. We can see that a higher feature size does not necessarily imply better performance. In fact, the best performance in mR@K and F1@K is reported by the smallest size 512. However, the latency and number of parameters scale according to the size, with a maximum of 281.7 ms for the biggest one 4096.
Table 6: Performance of the REACT model using different feature sizes for the edge and node representations.
Size mR@K R@K F1@K Latency Params
512 19.7 30.2 23.9 32.5 42.7M
1024 14.6 30.6 19.7 46.6 71.7M
2048 15.5 29.4 20.3 99.2 168.5M
4096 14.1 30.8 19.3 281.7 515.9M
@@ Acknowledgments
This work has been supported by the Brittany Region.
Supplementary Material
Appendix A Datasets
A.1 VG150 Annotations Biases
The VG150 dataset has become the main benchmark for Scene Graph Generation. This dataset has been extracted from original Visual Genome (VG) annotations. In their work, Xu et al. [41] proposed to select the top 150 objects and top 50 predicate classes from VG to compose the VG150 dataset. However, in this process, no manual refinement has been performed to verify the coherence of the data. The original Visual Genome annotations are known to be very noisy [45] with a lot of interchangeable classes. In Tab. 7 we display some of the classes of the VG150 dataset. For every column, we can see that all classes from subsequent rows can be used interchangeably with the label from the first row. This leads to strong confusion of (1) the object detector and (2) the relation predictor.
Objects Predicates
person animal vehicle wears on
girl dog car wearing on top of
woman bear truck above
man cat train
Table 7: Confused classes of VG150.
In Figure 5 we display the confusion matrix for the YOLOV8-medium model trained on VG150. We can clearly observe confusion between the class ”man”, ”boy” and ”person” or another confusion between ”airplane” and ”plane” which demonstrates the low quality of annotations. As a result, we decided not to compare our approach to the VG150 dataset. We still report the results of the REACT model on the VG150 dataset to compare with approaches that cannot be retrained on more qualitative datasets.
Figure 5: Normalized confusion matrix for the YOLOV8-m model on VG150. For convenience, we only display values above 0.1.
A.2 Other Datasets
In addition, we choose not to compare our approach on the GQA [8] and OpenImage [18] datasets. The rationale is as follows: the GQA dataset is also a split of Visual Genome and thus contains the same biases as VG150 [42]. In addition, GQA contains 301 predicate classes with a lot of them being ambiguous or that can be used interchangeably (such as ”standing” and ”standing on”, etc…). The OpenImage dataset contains only very sparse annotations (an average of 2.7 relations per image) which is not tailored for SGG. In addition, it includes a lot of noisy object classes with only a few samples each.
VG150, OpenImage, and GQA are all heavily long-tailed datasets. Our work does not focus on long-tail learning [35], in consequence, we report our main results on the PSG and IndoorVG datasets which are less biased toward tail classes. In addition, these datasets have more qualitative annotations with no spurious or ambiguous classes. For the IndoorVG dataset, we used the V4 version provided in the authors’ codebase. In Table 8 you can find a comparison of each dataset.
Table 8: Comparison of different Scene Graph datasets, # Rel. displays the total number of relationships and Rel/Img the average number of relations per image.
Dataset Images Obj. Pred. # Rel. Rel/Img
VG150 [17] 108,073 150 50 622,705 5.76
OpenImage [18] 133,503 601 30 367,914 2.75
GQA [8] 85,638 200 310 471,614 5.51
PSG [42] 48,749 133 56 278,356 5.71
IndoorVG [28] 14,674 83 34 112,804 7.69
Appendix B YOLO Versions Comparison
Table 9: Performance of the REACT model using different object detectors with the DTS architecture. We used the medium version of each detector.
Model mAP50 mR@K R@K F1@K Lat. Params
YOLOV8 53.1 19.7 30.2 23.9 32.5 42.7M
YOLOV9 53.3 16.3 29.8 21.0 46.5 39.7M
YOLOV10 54.5 16.9 33.0 22.3 55.3 33.4M
In the following, we compare the performance of our REACT model with different YOLO models for feature extraction and object detection. Similar to YOLOV8, we trained the YOLOV9-medium [39] and YOLOV10-medium [38] models on the PSG dataset [42] with batch size 8 for 100 epochs. Then, the REACT model has been trained for 20 epochs using the same hyperparameters as for YOLOV8. Results are displayed in Table 9. We observe no significant improvements in mR@K or R@K by using YOLOV9 or V10 in comparison to V8. The performance of YOLOV9 and YOLOV10 in object detection is superior to YOLOV8, by 0.2 and 1.4 points in mAP, respectively. Interestingly, the overall latency is worse for YOLOV10 and YOLOV9 than by using YOLOV8, which contradicts findings by the respective authors [39, 38]. For YOLOV9, we believe this could come from the size of the feature maps extracted from the backbone which is different than YOLOV8 (YOLOV9m uses a GELAN architecture with an output size of 240, whereas YOLOV8 output size is 192). Higher feature sizes will lead to more computation in the following stages as the subject and object visual representation will be based on these features.
Appendix C Detailed Latency
In the following, we detail the latency of each component of the Two-Stage and Decoupled Two-Stage architectures. Before feeding images to the model, some pre-processing is required (mainly image Transforms). To generate the final predictions, post-process is applied to convert class labels to names and rescale bounding boxes to the original image size. In Table 10 we display these values as well as a detail of the latency of the object detector (feature extraction + object detection) and the relation predictor. In addition, we display the minimum and maximum values obtained on the PSG dataset test set. Very low values account for no bounding box detected and high values for the maximum of 100 bounding boxes detected.
We observe significant differences in latency for the relation prediction stage while object detection stays relatively similar for all models. The difference between max and min values for DTS is 7.91x whereas for TS models it is 12.67x. In addition to a more lightweight backbone, this difference demonstrates that our DTS method is able to remove computation from the relation prediction head. The significant range between minimum and maximum latency is an important parameter that needs to be accounted for deployment. Our REACT model has the lowest difference between min and max values, with 2.85x, showing that our relation head is the most efficient for real-time constraints.
Table 10: Details of the latency for each model. The Min and Max correspond to the min and max values of the full forward pass of the model (without pre and post-process).
B Model Latency (ms)
Pre-process Obj. Det. Rel. Pred. Post-process Min Max
DTS REACT 0.60 16.08 15.70 0.53 19.40 55.36
PE-NET 0.60 17.18 81.42 0.78 25.14 199.54
GPS-NET 0.64 15.31 22.06 0.57 17.17 105.30
Neural-Motifs 0.60 14.82 19.68 0.57 14.00 115.01
VCTree 0.60 16.20 416.47 0.61 63.89 591.94
Transformer 0.57 14.94 17.31 0.51 17.33 138.74
TS PE-NET 0.97 91.35 287.63 0.62 118.09 1625.90
GPS-NET 0.95 86.10 223.75 0.55 105.83 1325.97
Neural-Motifs 1.01 86.19 224.96 0.53 113.87 1373.58
VCTree 0.99 82.10 320.77 0.51 127.54 1583.71
Transformer 0.94 85.20 201.00 0.53 105.96 1308.07
Figure 6: Comparison of Two-Stage and Decoupled Two-Stage architectures with our DCS strategy for inference.
Appendix D Dynamic Candidate Selection
D.1 Rationale
In the Object Detection task, a confidence threshold is traditionally selected at inference time to control the quality of output bounding boxes. The paradigm of SGG leverages an Object Detection component, however, applying a set threshold to control the number of predicted proposals before the relation prediction stage is not desirable. First, the Two-Stage approaches rely on a separate classifier to decode the box labels after the relation prediction stage. The performance of this classifier will impact the final score of the predicted proposals. Second, setting a hard threshold for object detection in our DTS approach will significantly affect the performance concerning relation prediction as the metrics used are not based on precision but rather recall (R@K and mR@K). This will then be unfair when compared with the two-stage approaches. As a result, controlling the number of input proposals by using a saturation threshold for all models (i.e. our DCS strategy) seems more promising. During inference in the wild, a threshold can be applied to the relation scores to only output high-confidence relations and corresponding box coordinates.
D.2 Models Comparison
In this section, we present an in-depth analysis of the DCS strategy. In addition to Decoupled Two-Stage models, we ran evaluation on Two-Stage architectures with Faster-RCNN and DCS on the PSG dataset, results are displayed in Figure 6. We plot the optimum value found for each model. We observe that the optimum for Two-Stage models is significantly higher than for Decoupled Two-Stage models. The average optimum for TS is 74, while for DTS it is 37. This difference can be explained by the low quality of the bounding box predicted by the Faster-RCNN in Two-Stage architectures.
Appendix E Training Details
Table 11: Hyperparameters used for the training of Two-Stage and Decoupled Two-Stage models.
Parameter Value
Optimizer SGD
Learning Rate 0.01
Momentum 0.9
Batch Size 8
Seed 42
Epochs 20
All models are trained using torch.use_deterministic_algorithms as well as a set seed for all random variables. In Table 11 we display the hyperparameters used for all training of Two-Stage and Decoupled Two-Stage models. These values are taken from previous work [34, 44, 46, 35, 26]. We trained the models for 20 epochs using early stopping (grace period=3). For each model, we evaluate the best epoch on the validation set (metric used: F1@K).

### v3

@@ B Model mR@K R@K F1@K mAP50
SQUAT [Jung et al.(2023)Jung, Kim, Kim, and Cho] 15.3 26.7 19.4 -
BGNN [Li et al.(2021)Li, Zhang, Wan, and He] 11.6 33.4 17.3 29.0
@@
OS EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] 7.8 29.3 12.4 30.8
SGTR [Li et al.(2022a)Li, Zhang, and He] 13.6 26.5 18.0 25.4
SS-RCNN [Teng and Wang(2022b)] 9.5 36.0 15.0 23.8
RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn] 8.0 24.3 12.0 26.4
Table 4: REACT compared to previous work on VG150 [Xu et al.(2017)Xu, Zhu, Choy, and Fei-Fei, Krishna et al.(2017)Krishna, Zhu, Groth, Johnson, Hata, Kravitz, Chen, Kalantidis, Li, Shamma, Bernstein, and Fei-Fei] (reported results). mR@K and R@K are the average for $k=[50,100]$ as some approaches do not report $k=20$.
@@
By using YOLOV8, we improved the mAP by 54.37% on the PSG dataset (Table˜2) and 69.92% on IndoorVG (Table˜4), compared to TS approaches with Faster-RCNN. Compared to OS approaches (EGTR [Im et al.(2024a)Im, Nam, Park, Lee, and Park] and RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn]), mAP is improved by 120% on average, which demonstrates the inefficiency of such approaches for the OD task. For a fair comparison, we also present the performance of the REACT model on the VG150 dataset, see Table˜4. The performance of REACT in mAP is better by one point than the best approach EGTR. We believe that the low quality of the bounding box annotations and the ambiguous classes in VG150 are confusing the YOLOV8 detector, which leads to only a small improvement against Faster-RCNN. On the PSG dataset, we experienced an average improvement of 58.47% in F1@k by using our DTS method compared to the same approaches with Faster-RCNN. In addition, our REACT model outperforms the PE-NET model with DTS by a small margin (23.9 versus 23.2 in F1@K). We observe that the gain in F1@k is not very strongly correlated to the gain in mAP between the Two-Stage and DTS approaches, as we measured a correlation coefficient of 0.75 between the two. On the IndoorVG dataset, we observe an improvement of 38.76% in F1@K with the DTS approach. Our REACT model also outperforms PE-NET (DTS) by a strong margin (22.8 versus 20.7 in F1@K). We observe that REACT outperforms some Two-Stage approaches (Motifs, VCTree, and BGNN) for RelPred on VG150 (see Table˜4) but is not competitive with more recent approaches (PE-NET and SQUAT). We believe this is due to noisy feature representation. Because REACT has fewer parameters, it is more sensitive to noisy data. Regarding latency, we demonstrated an improvement of 84.99% using YOLOV8 instead of Faster-RCNN, which is a considerable gap. Regarding the number of parameters, the DTS approach uses on average 77.47% less parameters than the baseline. Our introduced REACT model is the fastest one, with a latency of 32.5ms. By applying our DCS strategy, we gain an average of 66.5% latency on all models (compared to the YOLOV8m-DTS versions) with a relative average loss of 1% in F1@K. The EGTR model stays the best overall in terms of parameter count by a slow margin (42.5M versus 43.3M for REACT). By comparing REACT to the fastest model benchmarked, RelTR [Cong et al.(2022b)Cong, Yang, and Rosenhahn], REACT is 2.7x faster with +10+10 points on average in F1@K and +24.9+24.9 points on average in mAP50 for the 3 datasets. By taking into account OD accuracy, RelPred accuracy, and latency, our REACT model is the best compromise. Even for the VG150 dataset, REACT is a better choice because no other two-stage approaches are close to real-time inference. On the other hand, when evaluated fairly, one-stage approaches are clearly behind in terms of mAP and F1@K on all datasets, and are then not a good choice in comparison to REACT.
@@
Latency and Number of Proposals. As explained before, the second stage of SGG takes as input a fixed number of $n$ proposals as node candidates. There could be a maximum of n(n1)n*(n-1) possible pairs in the graph, thus when doing matching the computational complexity is supposed to scale accordingly. To demonstrate this hypothesis, we evaluated the performance of the REACT model in both latency and accuracy for different numbers of input proposals $k=\{10,150\}$ with DTS-YOLOV8 and with classical Faster-RCNN. For all experiments, we ranked proposals by confidence of the backbone and selected the top $k$.
@@
Figure 3: Left: Latency for the REACT model using a different number of proposals per image, with batch size 1. Right: Average F1@k for the REACT model with DTS-YOLO.
@@
The results of these experiments are shown in fig.˜3 (top), where we observe a plateau phenomenon for REACT with Faster-RCNN above 100 proposals. This behavior does not appear when using DTS+YOLOV8. By using 150 proposals with DTS+YOLOV8, the latency was similar to 10 proposals with Faster-RCNN. The evaluation of the REACT model under different sampling conditions with our DCS strategy is represented in fig.˜3 (bottom). In this figure, we can see that the optimal value was found at 23.89 for $k=42$, just before that the performance starts to be saturated.
@@
Features Fusion. During context learning of the REACT model, and more generally in a lot of approaches in SGG [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang, Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao, Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song], different feature sources are used to model the edge representation. First, union and spatial features are combined and then merged with the addition of the subject and object node features. The node features are created using visual and textual features obtained for each proposal. In our REACT model, we removed the dependency on the union features (see Figure˜2(d) (i)). In the following, we justify our choice by analyzing the impact of each feature type on learning. The results from these experiments are displayed in Table˜5. The baseline REACT model is the second row. With union features (first row) we observe a significant decrease in F1@K as well as latency. We hypothesize that union features are too noisy and confuse the model. This is consistent with previous work which highlights the importance of spatial features for the task [Jin et al.(2023b)Jin, Guo, Meng, Zhu, Xi, Wang, Mu, and Song]. In the last two rows, we removed all dependencies on contextual features (union and spatial) and kept only the concatenation of the subject and object node as initial edge features. We observe only a slight decrease by keeping only text features (last row). In fact, the model can retain up to 71% of the optimal performance by learning only from the textual GLoVe features. This suggests a trade off between text and visual features in the task.
@@
Table 5: Feature sources experiments. V and T are visual and textual features for <subject,object><subject,object>, and S and U are Spatial and Union features, respectively.
@@ Acknowledgments
This work has been supported by the Brittany Region and the Knut and Alice Wallenberg foundation and the Wallenberg AI, Autonomous Systems and Software Program (WASP).